### v1

# Enhanced DeepONet for 1-D consolidation operator learning: an architectural investigation

Abstract
Deep Operator Networks (DeepONets) have emerged as a powerful surrogate modeling framework for learning solution operators in PDE-governed systems. While their use is expanding across engineering disciplines, applications in geotechnical engineering remain limited. This study systematically evaluates several DeepONet architectures for the one-dimensional consolidation problem. We initially consider three architectures: a standard DeepONet with the coefficient of consolidation embedded in the branch net (Models 1 and 2), and a physics-inspired architecture with the coefficient embedded in the trunk net (Model 3). Results show that Model 3 outperforms the standard configurations (Models 1 and 2) but still has limitations when the target solution (excess pore pressures) exhibits significant variation. To overcome this limitation, we propose a Trunknet Fourier feature-enhanced DeepONet (Model 4) that addresses the identified limitations by capturing rapidly varying functions. All proposed architectures achieve speedups ranging from 1.5 to 100 times over traditional explicit and implicit solvers, with Model 4 being the most efficient. Larger computational savings are expected for more complex systems than the explored 1D case, which is promising. Overall, the study highlights the potential of DeepONets to enable efficient, generalizable surrogate modeling in geotechnical applications, advancing the integration of scientific machine learning in geotechnics, which is at an early stage.
keywords:
Scientific machine learning , Neural Operator , DeepONet , Terzaghi’s Consolidation PDE , Surrogate Modeling , Fourier Feature Embedding ,
[…]
[label1] organization=School of Civil and Environmental Engineering, Atlanta, Georgia Institute of Technology, city=Atlanta, postcode=30332, state=GA, country=USA
1 Introduction
The modeling and simulation of geotechnical problems typically rely on solving partial differential equations (PDEs) or ordinary differential equations (ODEs) using numerical methods such as the Finite Element Method (FEM) or the Finite Difference Method (FDM). While these approaches are well-established, they often incur significant computational costs.
Recent advances in scientific machine learning (Faroughi et al., 2022) have introduced computationally efficient deep learning-based surrogate modeling frameworks (Huang et al., 2025) to approximate the solutions of PDEs obtained using numerical solvers. Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019; Cuomo et al., 2022) constitute one of these emerging frameworks. PINNs approximate PDE solutions by enforcing governing equations as soft constraints within the loss function during neural network training. Although PINNs have been widely explored in various engineering domains, their application in geotechnical engineering remains relatively limited, with recent studies focusing on assessing the potential of PINN. For example, Bekele (2021) evaluated the PINN potential for forward and inverse problems in soil consolidation. Wang et al. (2024) evaluated PINN’s performance in the context of experimental consolidation data. Gao et al. (2023); Yang et al. (2025) applied PINN to seepage analysis and soil-water characteristic curves. Liu et al. (2025) assessed PINN’s potential in seismic site response analysis, leveraging Fourier feature embedding.
Although PINNs have shown promise for solving PDEs in many engineering disciplines, recent studies (Grossmann et al., 2024; Bueno et al., 2025) indicate that they are not necessarily more effective PDE solvers compared to traditional numerical methods. PINNs are designed to approximate solutions for a fixed set of PDE coefficients and boundary conditions, which means that they require retraining if those conditions change (Faroughi et al., 2022). This retraining process is computationally expensive (Sharma and Shankar, 2022), which poses challenges in the use of PINNs in applications with variable system configurations. Additionally, PINNs require an explicit formulation of governing equations, which limits their applicability in domains where physical behavior is described by empirical or phenomenological models rather than well-defined PDEs.
Neural operators (Kovachki et al., 2023; Li et al., 2020a; Lu et al., 2021; Li et al., 2020b) have emerged as an alternative in scientific machine learning that addresses some of the PINN’s limitations. Neural operators are a new type of deep learning framework designed to learn operators that map between infinite-dimensional function spaces. Unlike conventional deep learning methods that approximate solutions for specific inputs, neural operators learn the underlying solution operator that maps input functions (e.g., material property fields, boundary conditions, loadings) to output solution fields. This operator-based learning enables neural operators to generalize across varying initial and boundary conditions without retraining. Additionally, neural operators do not require explicit knowledge of governing equations, which makes them particularly well-suited for complex physical phenomena where a closed-form PDE is not defined. Compared to conventional numerical methods or PINNs, which need to solve the underlying PDEs for each new set of conditions (Koric et al., 2024; Jiang et al., 2024; Faroughi et al., 2022), neural operators offer significantly more efficient inference based on learned function mapping. These advantages make them particularly well-suited for surrogate models in engineering applications with complex operating conditions.
Among various neural operator architectures, Deep Operator Networks (DeepONet), proposed by Lu et al. (2021), have gained substantial attention due to their formulation. DeepONet is based on the universal approximation theorem for operators (Chen and Chen, 1995), which establishes its ability to approximate nonlinear solution operators of PDEs given sufficient network complexity and data coverage. This framework allows DeepONet to efficiently learn generalizable operators (Lu et al., 2022).
DeepONet can achieve greater data efficiency, faster training convergence, and improved accuracy (Lu et al., 2021) compared to conventional architectures, such as feed-forward networks (FNNs), sequence-to-sequence models (seq2seq), and residual networks (ResNets) in operator learning tasks. Furthermore, the architecture of DeepONet offers interpretability through its explicit separation of functional roles: the branch net learns to encode the input function, representing the coefficients of a latent basis, while the trunk net learns to output corresponding basis functions over the spatio-temporal domain, which are then combined with these coefficients to form the solution. This design mirrors the structure of functional mappings in infinite-dimensional spaces, where solution operators can be viewed as weighted combinations of basis functions parameterized by input functions (Naylor and Sell, 1982). This formulation is in alignment with the mathematical framework of operator learning (Kovachki et al., 2023).
DeepONet has been applied in various fields, such as fluid mechanics (Kumar et al., 2024; Bai et al., 2024), material science (Yin et al., 2022), structural mechanics (He et al., 2023; Ahmed et al., 2024), and thermodynamics (Koric and Abueidda, 2023; Kushwaha et al., 2024); these efforts have highlighted DeepONet’s potential in modeling complex physical processes. However, the potential of DeepONet in geotechnical engineering remains largely unexplored. This study aims to introduce DeepONet as a surrogate modeling approach for geotechnical problems, focusing on the one-dimensional (1D) consolidation phenomenon, one of the most prevalent geotechnical engineering problems. The motivation for exploring the consolidation problem is that it can serve as a benchmark to demonstrate DeepONet’s capabilities to the geotechnical community, while also highlighting the role of the different architectural adaptations explored in this study.
More specifically, this study presents a systematic investigation into DeepONet architectural choices tailored for the consolidation phenomenon. We first explain the standard architecture of DeepONet, and then explore three model learning architectures to get initial insights. The first two models follow DeepONet variants investigated previously (Tan and Chen, 2022; Jin et al., 2022), and the third model adopts a non-traditional architectural choice, with its better accuracy supported by the analytical solution structure of the consolidation problem. Insights and challenges identified from the various explored architectures are shared and utilized to propose a DeepONet architecture enhanced with Trunknet Fourier feature embedding, which demonstrates superior performance. Overall, this study contributes to the integration of advanced scientific machine learning into geotechnical applications.
2 Deep operator networks(DeepONet)
DeepONet approximates mappings between infinite-dimensional function spaces. This capability positions DeepONet as a powerful framework for operator learning. Formally, consider an operator $\mathcal{G}$ defined as:
$\mathcal{G}:\mathcal{U}\ni{u}\mapsto\mathcal{S}\ni s$ (1)
where $\mathcal{U}$ and $\mathcal{S}$ represent infinite-dimensional function spaces. Specifically, the operator $\mathcal{G}$ takes an input function $u\in\mathcal{U}$ and maps it to an output function $s\in\mathcal{S}$. DeepONet aims to approximate $\mathcal{G}$ through a neural network model $\mathcal{G}_{\Theta}$, such that:
$\mathcal{G}_{\Theta}(u)(y)\approx s(y),\quad\text{for all }y\in D,$ (2)
Here, $y$ denotes a point within the spatial or temporal domain $D$ of interest. The input function $u$ is sampled at a fixed set of $m$ locations, ${\{x_{\ell}\}_{\ell=1}^{m}}$, referred to as sensor locations. $\Theta$ is a set of learnable parameters of the network.
Structurally, DeepONet comprises two subnetworks: the branch network (branch net) and the trunk network (trunk net) (see Figure 1). The branch net encodes the input function $u$, producing a latent feature representation $\boldsymbol{b}(u)\in\mathbb{R}^{q}$. The trunk net encodes the evaluation location $y$, generating another latent feature representation $\boldsymbol{t}(y)\in\mathbb{R}^{q}$. The final operator evaluation at a specific location $y$ is obtained via the inner product of these two latent representations as follows:
$\mathcal{G}_{\Theta}(u)(y)=\sum_{k=1}^{q}b_{k}(u(x_{1}),u(x_{2}),...,u(x_{m}))t_{k}(y)$ (3)
where $b_{k}(u)$ and $t_{k}(y)$ denote the $k$-th component of the branch and trunk net outputs, respectively. The trunk and branch net are modeled with neural networks. The parameters of the neural networks comprise $\Theta$. The output of the trunk net can be interpreted as a basis for representing the output function, and the output of the branch net as the corresponding coefficients. DeepONet learns an optimal set of basis functions directly from data, allowing it to adapt to complex solution manifolds.
DeepONet’s learned basis and corresponding coefficients eliminate the need for manually designed basis functions, which are typically required in traditional numerical methods. For example, the finite element method uses local basis functions defined over a mesh, while spectral methods rely on global functions like sine and cosine series (e.g., Fourier series) chosen in advance (Williams et al., 2024). This flexibility allows DeepONet to effectively capture complex, nonlinear operator mappings in high-dimensional PDEs.
Figure 1: A standard DeepONet architecture.
3 Consolidation and associated PDEs
Consolidation describes how excess pore water pressure (PWP) dissipates from saturated soils under load over time, causing gradual ground settlement, particularly in clayey soils. The governing PDE for 1D consolidation under vertical drainage (Equation 4) is derived from Darcy’s law for fluid flow through porous media and the effective stress principle.
$\frac{\partial u}{\partial t}=C_{v}\frac{\partial^{2}u}{\partial z^{2}}$ (4)
where $u(z,t)$ is the excess PWP at depth $z$ and time $t$, and $C_{v}$ is the coefficient of consolidation, a parameter that combines the effects of soil permeability and compressibility.
Equation 4 represents a diffusion-type process: pore pressure gradients drive vertical flow through the soil matrix, and as water escapes, the excess pore pressure gradually diminishes. This dissipation results in an increase in effective stress and ultimately leads to ground settlement.
A common drainage condition to solve equation Equation 4, which is used in this study, considers that drainage occurs only through the top surface of the soil, while the bottom boundary is considered impermeable. This leads to a combination of boundary conditions: a Dirichlet condition at the top $u(0,t)=0$, and a Neumann condition at the bottom $\frac{\partial u}{\partial z}\bigg{|}_{z=H}=0$, which enforces a zero hydraulic gradient. The analytical solution to Equation 4 is given by:
$u(z,t)=\sum_{m=0}^{\infty}\left[\frac{2u_{0}}{M}\sin\left(\frac{Mz}{H_{\mathrm{dr}}}\right)\right]e^{-M^{2}T_{v}}$ (5)
Where $M=\frac{\pi}{2}(2m+1)$ under single drainage boundary condition, $u_{0}$ is the initial excess pore pressure, $H_{dr}$ is the drainage path length, and $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}$ is the dimensionless time factor. This analytical form reveals that the solution is composed of spatial sine functions modulated by exponential decay terms dependent on $C_{v}$ and time $t$. This insight informs one of the DeepONet architecture choices that is later explored.
4 Data generation and training
We use the finite difference method to numerically solve Equation 4. The solution obtained is used to train the DeepONet and to test the performance. The finite difference method first discretizes the spatial domain uniformly into $N_{z}$ nodes separated by a spacing $\Delta z$. It then approximates the spatial second derivative at each internal node $i$ ($21\leq i\leq N_{z}-$ 2), using a second-order central finite-difference scheme (Equation 6).
$\frac{\partial^{2}u}{\partial z^{2}}\bigg{|}_{z=z_{i}}\approx\frac{u_{i+1}-2u_{i}+u_{i-1}}{\Delta z^{2}}.$ (6)
This discretization converts the PDE into a semi-discrete system of ODE (Equation 7).
$\frac{du_{i}}{dt}=C_{v}\frac{u_{i+1}-2u_{i}+u_{i-1}}{\Delta z^{2}},\quad 1\leq i\leq N_{z}-2.$ (7)
For the single drainage boundary condition, the Dirichlet condition $u_{0}=0$ is applied at the top ($z=0$). At the impermeable bottom boundary ($z=H$, node $N_{z}-1$), the Neumann condition ($\frac{\partial u}{\partial z}|_{z=H}=0$) is applied by considering a ghost node $u_{N_{z}}$ and setting $u_{N_{z}}=u_{N_{z}-1}$. The ODE for the bottom boundary node $N_{z}-1$ thus becomes:
$\frac{du_{N_{z}-1}}{dt}=C_{v}\frac{u_{N_{z}-1}-2u_{N_{z}-1}+u_{N_{z}-2}}{\Delta z^{2}}=C_{v}\frac{u_{N_{z}-2}-u_{N_{z}-1}}{\Delta z^{2}}.$ (8)
We employ both the implicit Backward Differentiation Formula (BDF) and the explicit Runge-Kutta (RK) methods to solve the discretized ODE system (see Equation 7 and Equation 8). BDF is used to generate data for DeepONet, while the RK method serves as a benchmark for comparative analysis of the computational performance of DeepONet and BDF. Both integrators are implemented using the solve_ivp function from the SciPy library, with details included in A.
When generating data, we consider both uniform and spatially varying initial excess PWP profiles $u(z,0)$ as the initial condition. For the uniform profile, we randomly sample constant functions between 10 kPa and 20 kPa. For the spatially varying (non-uniform) profiles, we sample functions from a Gaussian Random Field (GRF), denoted as $u(z,0)\sim\mathcal{G}(m,C(z_{1},z_{2}))$. The spatial correlation of this GRF is defined by a Gaussian covariance function:
$C(z_{1},z_{2})=\sigma^{2}\exp\left(-\frac{||z_{1}-z_{2}||^{2}}{l^{2}}\right)$ (9)
Here, $m$ is mean, $C(z_{1},z_{2})$ represents the covariance between two points at depths $z_{1}$ and $z_{2}$. The term $||z_{1}-z_{2}||$ is the Euclidean distance between these points. The parameter $l$ is the correlation length (or length-scale), which controls the smoothness of the sampled functions; a larger $l$ generates smoother profiles. The term $\sigma^{2}$ is the variance of the field, controlling the overall magnitude of the covariance. We use a $\sigma^{2}$ value of 1000 $kPa^{2}$, a correlation length of 0.5, and randomly select $m$ from 10 kPa to 20 kPa. The coefficient of consolidation, $C_{v}$, is sampled uniformly from $[0.3,1.0]$ m2/year. The spatial domain spans the normalized depth $\frac{z}{H_{dr}}\in[0,1]$ and normalized time $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}\in[0,2]$.
In the DeepONet training, the model parameter set $\Theta$ in $\mathcal{G}_{\Theta}$ is optimized by minimizing the mean squared error between the predicted solution and the true solution at $y$ as shown in Equation 10.
$\mathcal{L}(\Theta)$ $=\frac{1}{NP}\sum_{i=1}^{N}\sum_{j=1}^{P}\left|\mathcal{G}_{\Theta}(u^{(i)})(y_{j}^{(i)})-\mathcal{G}(u^{(i)})(y_{j}^{(i)})\right|^{2}$ (10)
$=\frac{1}{NP}\sum_{i=1}^{N}\sum_{j=1}^{P}\left|\sum_{k=1}^{q}b_{k}(u^{(i)}(x_{1}),\dots,u^{(i)}(x_{m}))t_{k}(y_{j}^{(i)})-\mathcal{G}(u^{(i)})(y_{j}^{(i)})\right|^{2}$
where $\{u^{(i)}\}_{i=1}^{N}$ denotes $N$ separate input functions sampled from $\mathcal{U}$. For each $u^{(i)}$, $\{y_{j}^{(i)}\}_{j=1}^{P}$ are $P$ locations in the domain of $\mathcal{G}(u^{(i)})$, and $\mathcal{G}(u^{(i)})(y_{j}^{(i)})$ is the corresponding output data evaluated at $y_{j}^{(i)}$. Contrary to the fixed sensor locations of $\{x_{\ell}\}_{\ell=1}^{m}$, the locations of $\{y_{j}^{(i)}\}_{j=1}^{P}$ may vary for different $i$, thus allowing to construct a flexible and continuous representation of the output functions $s\in\mathcal{S}$.
We train and validate DeepONet using N=40,000 and N=5,000 input functions, respectively. Each input function represents an initial excess PWP and is discretized at m=100 fixed, equally spaced sensor locations, denoted by $\{x_{\ell}\}_{\ell=1}^{100}$ (see the left side figure in Figure 2a). That is, for each input function $u^{(i)}$, the branch network receives the vector $\left[u^{(i)}(x_{1}),u^{(i)}(x_{2}),\dots,u^{(i)}(x_{100})\right]$ as input. The corresponding output functions (i.e., solutions $s=\mathcal{G}(u)$) are sampled at P=100 randomly selected evaluation points $y=(z,t)$ (see the red markers in the right side figure in Figure 2a), sampled independently for each input function. These locations are denoted $\{y_{j}^{(i)}\}_{j=1}^{100}$, resulting in a total of $40,000\times 100$ training examples of the form $\left(u^{(i)},y_{j}^{(i)},\mathcal{G}(u^{(i)})(y_{j}^{(i)})\right)$. The validation set is constructed in the same manner, with P=100 output evaluations per input, resulting in $5,000\times 100$ training examples. For testing, we assess the generalization capability of the learned operator $\mathcal{G}_{\Theta}$ on 500 unseen input functions during training. Each is evaluated over a dense $100\times 100$ uniformly spaced grid in the output domain (see the right side figure in Figure 2b), yielding P=10,000 output locations per function to visualize full solution fields. All input and output data are standardized using the mean and standard deviation computed from the training set.
Figure 2: (a) Training and Validation Data Sampling Method. Input functions (initial excess pore water pressure (PWP)) are sampled at $m=100$ fixed equally spaced locations, denoted by $\{x_{\ell}\}_{\ell=1}^{100}$ (see left side figure). For each given input function $u^{(i)}$, the corresponding solution is sampled at $P=100$ randomly chosen output evaluation points, denoted by $\{y_{j}^{(i)}\}_{j=1}^{100}$ (the red markers on the right side figure). This results in 100 training examples of the form $(u^{(i)},y_{j}^{(i)},\mathcal{G}(u^{(i)})(y_{j}^{(i)}))$ for each input function. For the training data, we generate $N=40,000$ input functions, forming a total of $40,000\times 100$ training examples. For validation, we generate $N=5,000$ input functions, forming $5,000\times 100$ validation examples. (b) Test Method. Input functions unseen during training are sampled at $m=100$ fixed equally spaced locations, $\{x_{\ell}\}_{\ell=1}^{100}$ (see left side figure). For each given input function $u^{(i)}$, we evaluate the learned solution operator $\mathcal{G}_{\Theta}(u^{(i)})(y)$ at a dense $100\times 100$ uniformly spaced grid of $y$ points across the output domain (see right side figure).
5 Model architectures
We explore various model architectures, each reflecting a modeling assumption about the interaction between the PDE parameters, the input function, and the resulting solution. The motivation is to explore how the treatment and placement of governing parameters can affect performance. Figures 3, 4 and 5 shows the different architectures investigated, and described as follows:
  • 1.
    Model 1 (Figure 3): This model serves as our baseline and adheres to the standard DeepONet formulation introduced by Lu et al. (2021). The branch net processes the input function, which in this case consists of the discretized initial PWP profile, $u(z,0)$. The scalar coefficient of consolidation, $C_{v}$, is concatenated with the discretized $u(z,0)$ vector, forming a single input tensor for the branch net. This design philosophy treats $C_{v}$ as a global conditioning parameter. The implicit assumption is that the branch net should learn a combined representation of how the initial PWP distribution and the material property $C_{v}$ jointly influence the consolidation process. The trunk net takes the spatiotemporal coordinates $(z,t)$ as input.
  • 2.
    Model 2 (Figure 4): This architecture modifies the branch net to handle $u(z,0)$ and $C_{v}$ separately before integrating their influence. One branch net is dedicated to encoding the initial PWP function, $u(z,0)$, while a second, auxiliary branch net processes the coefficient of consolidation, $C_{v}$, generating distinct latent representations for each. These two latent representations are then concatenated and passed through a ”merge net” (another small neural network) to produce the final branch output vector. This approach allows the model to potentially learn more disentangled features for the initial condition and the PDE coefficient before their combined effect on the solution basis is determined. The trunk net architecture remains the same, taking $(z,t)$ as input.
  • 3.
    Model 3 (Figure 5): In Model 3, we propose feeding $C_{v}$ directly into the trunk net. The initial PWP function, $u(z,0)$, is processed by the branch net as before. However, the trunk net now receives an augmented input vector $(z,t,C_{v})$. The rationale is that the trunk net learns to approximate the basis functions of the solution operator. By providing $C_{v}$ directly to the trunk, it can generate basis functions dynamically modulated by this physical parameter across the spatiotemporal domain.
Figure 3: DeepONet with direct concatenation of $C_{v}$ to the branch net.
Figure 4: DeepONet with auxiliary networks for embedding $u(z,0)$ and $C_{v}$ and merging their representations.
Figure 5: DeepONet with trunk net modulation by $C_{v}$.
The performance of Models 1, 2, and 3 is discussed in Section 6, highlighting the challenges to predict excess PWP distribution during the early stages of consolidation, especially for non-uniform distributions across depth. This motivated the authors to explore an additional option. Specifically, an additional architecture option considers Model 3 enhanced with Fourier Feature Embedding (Tancik et al., 2020), which can help to learn rapidly varying functions (e.g., Wang et al. (2021), Liu et al. (2025)).
Fourier feature embedding maps low-dimensional inputs, such as spatial coordinate $z$ and time $t$ in our case, into a higher-dimensional space using sinusoidal functions of multiple frequencies. Formally, each input is transformed via:
$\gamma(\textbf{v})=\begin{bmatrix}\sin{(2\pi\mathbf{B}\textbf{v}})\\ \cos{(2\pi\mathbf{B}\textbf{v}})\end{bmatrix}^{T}$ (11)
where v is the input feature vector (trunk net input in this problem) of length $k$ and $\gamma(\textbf{v})$ is the embedded Fourier feature vector with a length of $2m$. $\mathbf{B}$ is a $m$ by $k$ matrix of frequencies, sampled from $\mathcal{N}(\mathbf{0},\sigma^{2}I)$. The tunable hyperparameters $m$ and $\sigma$ are pre-specified before the neural network training. This embedding introduces periodic features that enrich the input representation, allowing neural networks to approximate functions with high variation. Leveraging the Fourier feature embedding, we propose the following architecture:
  • 1.
    Model 4 (Figure 6): The DeepONet architecture is the same as Model 3; however, the Fourier feature embedding maps the input $(z,t,C_{v})$ into a higher-dimensional latent space before it is passed to the trunk net. This embedding transforms the original low-dimensional input into a richer representation using sinusoidal functions of varying frequencies, enabling the trunk net to learn functions with large variations.
Figure 6: DeepONet with trunk net modulation by $C_{v}$ and Fourier feature embedding
All the networks (i.e., branch, trunk, and merge net) in DeepONet models in this study employ fully connected multilayer perceptrons (MLPs). Each MLP comprises 6 layers with 30 neurons per layer. The output dimension of the branch and trunk net for the dot product is set to $q=50$ ( Equation 3). For Fourier feature embedding hyperparameters (Equation 11) in model 4, we use $m=50$ and $\sigma=1.0$.
6 Results
This section presents the performance evaluation of the four investigated DeepONet architectures in predicting excess PWP dissipation for the consolidation problem. We assess the models based on their accuracy under varying initial excess PWP $u_{0}$ and coefficient of consolidation $C_{v}$ conditions, and their computational efficiency.
Figure 7 shows the comparative performance of the four DeepONet architectures for a representative case in the test data with a non-uniform $u_{0}$ and a $C_{v}=0.5\ m^{2}/year$. A non-uniform $u_{0}$ is selected as it imposes fluctuating responses early in the consolidation process and enables a more rigorous assessment of the implemented architectures. Moreover, non-uniform $u_{0}$ profiles appear in geotechnical practice (Lovisa and Sivakugan, 2015), for instance, due to staged construction, uneven surface loading, or localized drainage.
The first row of Figure 7 shows the initial excess PWP function and the true solution. After this row, each of the figures corresponds to models 1 to 4. For each model, three subplots are presented: (1) the predicted spatiotemporal excess PWP field with MSE value in the figure title, (2) the corresponding absolute error with maximum error value in the title, and (3) a comparison of predicted and true pore pressure profiles at selected time snapshots.
All models capture the overall spatiotemporal excess PWP dissipation trends; however, the accuracy varies depending on the models and the region. Model 1, the standard DeepONet with $C_{v}$ concatenated to the initial excess PWP in the branch net, yielded a MSE of $9.38\times 10^{-4}\ Pa^{2}$ and a maximum absolute error of $1.40\times 10^{3}\ Pa$. Model 2, which processes $C_{v}$ in a separate branch net before merging, exhibits a higher MSE of $1.14\times 10^{-3}\ Pa^{2}$ and a maximum absolute error of $1.31\times 10^{3}\ Pa$. While both models capture the overall spatiotemporal trend, they display noticeable deviations from the true solution and larger errors during the early stages of consolidation, particularly in regions with the steepest pore pressure gradients.
Model 3, where we incorporate $C_{v}$ as an input to the trunk net, shows a significant improvement in prediction accuracy. This architecture achieves an MSE of $1.17\times 10^{-4}\ Pa^{2}$ and a maximum absolute error of $8.69\times 10^{2}\ Pa$. The predictions from Model 3 align more closely with the true solution across the spatiotemporal domain, with substantially reduced absolute errors, especially in areas of rapid excess PWP change. However, slight elevations in error are still present at the very initial part of the consolidation process for this model.
Model 4, which augments Model 3 by applying Fourier feature embedding to the trunk net’s input $(z,t,C_{v})$, demonstrates the highest accuracy. It achieves the lowest MSE of $3.69\times 10^{-5}\ Pa^{2}$ and the smallest maximum absolute error of $2.97\times 10^{2}\ Pa$ for this case. Model 4 effectively minimizes the errors that the other models show, particularly reducing errors at the early stages of consolidation where excess PWP pressure changes are most pronounced.
In the time snapshot plots (rightmost subplots), early-stage prediction errors become more evident. Models 1 and 2 exhibit large discrepancies in the pore pressure distribution during the initial phase (from $T_{v}=0.0$ to $0.028$), not accurately capturing the rapid variation in pore pressure. In contrast, Models 3 and 4 more accurately reproduce the early-time profiles. However, Model 3 still shows a slight mismatch at $T_{v}=0.0$. This residual error is effectively mitigated by Model 4, which demonstrates improved accuracy at the earliest time step, owing to the enhanced trunk net input representational capacity introduced by the Fourier feature embedding.
Figure 7: Excess pore water pressure (PWP) dissipation evaluated by the different DeepONet architectures on test data at $C_{v}=0.5\ m^{2}/year$ with varying initial excess PWP over the depth. The figures in the first row show the initial excess PWP function and the true solution. The subsequent rows show the prediction from each model with the absolute error compared to the true solution, and the plots comparing the true and predicted excess PWP distributions over depth at selected times.
Figure 8 presents the model performances for a case in the test data with a uniformly distributed initial excess PWP with $C_{v}=0.7\ m^{2}/year$. The configuration of the figure is the same as Figure 7. In this simpler case, overall errors tend to be smaller than those from the non-uniform case shown in Figure 7. The concentration of errors near the early stage of consolidation is similarly evident as in the case shown in Figure 7.
Model 1 achieves an MSE of $4.29\times 10^{-5}\ Pa^{2}$, with moderate absolute errors predominantly concentrated near the early time steps. Model 2 again shows the weakest performance, recording the highest MSE of $1.44\times 10^{-4}\ Pa^{2}$ and the largest maximum absolute error of 955 Pa.
Model 3, with $C_{v}$ in the trunk net, again demonstrates a clear advantage, yielding an MSE of $1.70\times 10^{-5}\ Pa^{2}$ and a maximum absolute error of $1.55\times 10^{2}\ Pa$. Its predictions accurately track the time-dependent behavior with reduced absolute error. Model 4 achieves the highest accuracy by further improving upon Model 3 through the incorporation of Fourier feature embedding. Model 4 results in the lowest MSE of $1.51\times 10^{-5}\ Pa^{2}$ and the smallest maximum error of $107Pa$.
In the time snapshot plots (rightmost subplots), pressure distributions are mostly accurate for all models, including the early stages. However, we observe a few fluctuations at $T_{v}$ = 0.0 for model 1 and 2, and this fluctuation extends to $T_{v}=0.028$. Model 4 accurately captures all the early pressure distributions.
Figure 8: Excess pore water pressure (PWP) dissipation evaluated by the different DeepONet architectures on test data at $C_{v}=0.7$ with constant initial excess PWP over the depth. The figures in the first row show the initial excess PWP and the true solution. The subsequent rows show the prediction from each model with the absolute error compared to the true solution, and the plots comparing the true and predicted excess PWP distributions over depth at selected times.
Figure 9 shows the mean and standard deviation of MSE for each DeepONet architecture across training, validation, and test datasets, with a table showing the corresponding values. Models 1 and 2 consistently exhibit the highest MSEs and larger standard deviations. Model 2, featuring separate branch networks for the initial condition and $C_{v}$ performs the poorest, with a test MSE of 5921.96±5266.16 $Pa^{2}$ (mean ± 1 standard deviation). Model 3 significantly reduces the MSE across all datasets, achieving a test MSE of 740.62 ± 787.00 $Pa^{2}$, which is about 5 to 7 times lower than models 1 and 2. Model 4 shows the most accurate and consistent performance. The application of Fourier feature embedding to the trunk input $(z,t,C_{v})$ in Model 4 results in the lowest test MSE values, recording a test MSE of 508.96 ± 339.44 $Pa^{2}$, which further reduces the MSE of the model by about 1.5 times, along with markedly reduced variance compared to other models.
Figure 9: Mean and standard deviation of MSE for each DeepONet architecture across training, validation, and test datasets.
To further investigate the impact of Fourier feature embedding, we compare the worst-case predictions of the best-performing architectures: Model 3 and Model 4. These cases, shown in Figure 10, correspond to the samples with the highest mean squared error (MSE) for each respective model. Note that the color bars for the prediction and absolute error fields differ between the two models to reflect their distinct value ranges.
Model 4 demonstrates superior performance even in its worst-case scenario, with an MSE of 3.47e+03 $Pa^{2}$, approximately three times lower than the worst-case MSE of 9.87e+03 $Pa^{2}$ of model 3. A closer examination of the spatiotemporal absolute error distributions reveals key differences. It shows that model 3 exhibits highly localized error spikes, particularly at early times characterized by rapid excess PWP dissipation and high variability. The temporal comparison plots (rightmost column in each figure) for Model 3 show difficulty in matching the true pressure profiles during the initial phase ($T_{v}$=0.0 to 0.012). In contrast, Model 4 maintains a more uniformly distributed error profile with a generally reduced magnitude across the domain and achieves closer alignment with the true solution at all selected time snapshots, even in this worst-case prediction.
((a)) Model 3 ($C_{v}=0.5771\ m^{2}/year$)
((b)) Model 4 ($C_{v}=0.3016\ m^{2}/year$)
Figure 10: Worst case predictions for Model 3 and Model 4. Note that the figures do not share the same color bar scale for the prediction and error fields.
Figure 11 presents the training and validation loss curves for the four DeepONet architectures. All models show a consistent alignment between training and validation curves, indicating the generalization and effectiveness of DeepONet’s solution operator learning. However, each model achieves different loss levels. Model 1 and 2 demonstrate gradual loss reduction with mild fluctuations, but their convergence stabilizes at a relatively higher loss level. Model 2 converges at around 200 epochs, but this stabilization does not translate into strong predictive performance. Model 3 achieves notably lower losses and tends to keep decreasing even after 600 epochs, reflecting its superior performance. Model 4, while displaying some oscillations during early epochs, likely due to the increased complexity from Fourier feature embedding, ultimately attains the lowest overall losses, confirming its enhanced expressive capacity and improved accuracy.
Figure 11: Training and validation loss curves for DeepONet architectures.
Table 1 provides a comparison of the average computation times to generate a full spatiotemporal excess PWP solution field for the four DeepONet architectures and two classical numerical solvers (implicit BDF and explicit Runge-Kutta RK45). The reported times are averaged over 500 test cases, each corresponding to a distinct initial excess PWP profile (i.e., input function), and reflect the time to compute the entire solution field per case on an NVIDIA RTX 2000 Ada Laptop GPU. All DeepONet models demonstrate rapid prediction capabilities, with average execution times ranging from approximately 0.006 to 0.008 seconds per case. Model 4 is the fastest DeepONet architecture with a mean prediction time of 0.00619 s, followed by Model 3 (0.00677 s), Model 1 (0.00709 s), and Model 2 (0.00822 s).
Comparatively, the implicit BDF solver requires an average of 0.01118 s to produce a solution field under the same conditions used to evaluate DeepONet computation time, while the explicit RK45 solver is significantly slower, averaging 0.71304 s. Thus, the trained DeepONet models are approximately 1.5 to 2 times faster than the implicit numerical solver and about 100 times faster than the explicit numerical solver.
While the absolute time differences are marginal (order of milliseconds to sub-seconds) for 1-D consolidation (table 1), the results emphasize DeepONet’s potential as a computationally efficient surrogate model. This efficiency could be particularly significant when modeling large-scale or high-dimensional problems where conventional solvers may be computationally costly. Once trained, DeepONet enables rapid inference at any queried spatiotemporal regions without time-stepping across varying initial and material conditions and without retraining.
Table 1: Comparison prediction/execution time (in seconds) for DeepONet architectures and classical numerical solvers.
Method Mean (s) Std (s)
Model 1 (Standard) 0.00709 0.00281
Model 2 ($C_{v}$ Branch) 0.00822 0.00316
Model 3 ($C_{v}$ Trunk) 0.00677 0.00187
Model 4 ($C_{v}$ Trunk + Fourier) 0.00619 0.00166
Implicit solver (BDF) 0.01118 0.00095
Explicit solver (RK45) 0.71304 0.11409
7 Discussion: Performance of evaluated models and assessment of out-of-distribution data
A key finding in the result is the substantially improved prediction accuracy that Model 3 achieves, which takes the coefficient of consolidation, $C_{v}$, as an input to the trunk net rather than the branch net (i.e., models 1 and 2), which is the classical approach. This improved accuracy is supported by model 3’s architectural similarity to the analytical solution form of the consolidation equation (Equation 5). In Equation 5, $C_{v}$ is an intrinsic component of the dimensionless time factor, $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}$, which directly modulates the exponential decay terms (e.g., $e^{-M^{2}T_{v}}$). These exponential terms essentially act as the time-dependent basis functions in the solution. The trunk net in DeepONet is primarily responsible for learning and representing these basis functions of the solution operator. By providing $C_{v}$ directly as an input to the trunk net alongside the spatiotemporal coordinates $(z,t)$, the architecture better conditions the network to learn these $C_{v}$-dependent basis functions. This architectural alignment with the analytical solution allows Model 3 to more effectively capture the influence of $C_{v}$ on the spataio-temporal evolution of the excess PWP field. From a geotechnical perspective, $C_{v}$ encapsulates the combined effects of soil permeability and compressibility, which are fundamental in predicting the overall consolidation timeline. An architecture that directly learns this influence on the temporal basis functions, like Model 3 does, is therefore more likely to align with the physical principles of the consolidation process. Moreover, Fourier feature embedding effectively improves the accuracy of model 3 (i.e., model 4) (Figure 10), particularly in the early stages of consolidation. By mapping the low-dimensional inputs $(z,t,C_{v})$ into a higher-dimensional space via sinusoidal functions, Fourier feature embedding enriches the input representation and enables the network to more readily approximate functions with high-frequency content or sharp spatio-temporal patterns. This architectural refinement accounts for Model 4’s superior accuracy, especially in challenging scenarios with steep excess PWP gradients. These observations can guide the implementation of future DeepONet architectures in geotechnical problems with similar PDE-driven systems governed by system coefficients with steep functional variations.
We also evaluate the performance of the implemented DeepONet architecture on out-of-training distribution data. First, we assess the solution accuracy of Model 4 for values of $C_{v}$ outside the training range (Figure 12). For this, we create an extra test data set with $C_{v}$ randomly sampled from 0.1 to 1.5 $m^{2}/year$. Within the training range of $C_{v}$ (0.3 to 1.0 $m^{2}/year$, indicated by the shaded region), the model maintains low MSE, typically below $10^{3}\ Pa^{2}$. However, as $C_{v}$ goes into the out-of-distribution region, the MSE shows an increase. This behavior is not unexpected (e.g., (Lu et al., 2021; Ahmed et al., 2024)) because DeepONet is an effective parameter-aware interpolator rather than an extrapolator. This implies that the successful deployment of a DeepONet-based surrogate model requires a careful selection of PDE parameters in the training dataset, which, in turn, requires relevant domain knowledge.
Figure 12: Mean squared error evaluated on $C_{v}$s with out-of-training-distribution values for Model 4.
We further evaluate the solution accuracy of Model 4 on function spaces characterized by correlation lengths ranging from 0.2 to 0.8, extending beyond the training correlation length of 0.5. For each correlation length, we construct an additional test dataset comprising 10 realizations of Gaussian random fields as input functions. The lower panel in Figure 13 shows examples of these functions at the correlation lengths 0.2, 0.5, and 0.8. Consistent with prior work on DeepONet extrapolation (Lu et al., 2021), the prediction error increases when testing on functions with correlation lengths smaller than the 0.5 used in training. However, the prediction remains within a reasonable range; the mean MSE at a correlation length of 0.2 (approximately $6\times 10^{3}\ Pa^{2}$) is still lower than what we discussed in Figure 10(a), which exhibits an accurate solution except for the early times. Notably, the error does not increase for correlation lengths greater than 0.5, but decreases. This stability likely stems from the fact that GRFs with larger correlation lengths are smoother and more similar to the training function space, which spans 0.5 GRFs and uniform excess PWP functions. As a result, the model effectively interpolates between these GRFs and the constant functions. Along with Figure 12, the result in Figure 13 highlights the importance of careful DeepONet training data design based on domain knowledge.
Figure 13: Mean squared error evaluated on different function spaces with out-of-training-distribution correlation lengths (0.2 to 0.8) for Model 4. Vertical bars indicate one standard deviation. The bottom plots show several realizations of Gaussian random fields for correlation lengths of 0.2, 0.5, and 0.8, illustrating how the input function varies with different correlation length values.
8 Conclusion
This study demonstrates the performance of the DeepONet for learning solution operators in the 1-D consolidation problem, focusing on systematically investigating different DeepONet architectures. Our results reveal that feeding $C_{v}$ into the trunk net (model 3), rather than the more typical branch net (models 1 and 2), substantially improves prediction accuracy by 6 to 12 times lower test MSE. The improved accuracy is supported by the analytical form of the PDE, where $C_{v}$ modulates time-dependent behavior in the basis function. In addition, we further enhance the DeepONet architecture in Model 3 using Fourier Feature Embedding (Model 4). This further reduces the error by 1.5 times that of model 3, achieving the highest accuracy among all models. The architecture in Model 4 also effectively mitigates the high errors shown in highly varying, early-time excess PWP behaviors in the other model, due to its enhanced capacity to represent sharp spatio-temporal features from Fourier feature embedding. These findings provide guidance on architecture design for operator learning in similar PDE-driven systems governed by system coefficients.
The trained DeepONet models achieve a significant computational speedup compared to traditional numerical solvers used to generate training data and benchmarks. Although the absolute time differences are on the order of milliseconds for the 1D consolidation case, the relative speedup–ranging from 1.5 to 100 times faster–highlights DeepONet’s potential as an efficient surrogate model. This computational advantage could be particularly valuable in large-scale or high-dimensional geotechnical simulations, where conventional solvers incur high costs due to time-stepping and meshing overhead.
Our results also show that the trained DeepONet model performance diminishes when extrapolating to out-of-distribution values of $C_{v}$ and the function spaces for the initial excess PWP condition. This suggests that the deployment of a DeepONet-based surrogate model requires careful PDE parameter selection of a training dataset, which, in turn, requires relevant geotechnical expert knowledge.
Beyond these technical contributions, this work introduces DeepONet to the geotechnical engineering community as an effective surrogate modeling framework. Its ability to generalize across varying PDE parameters and initial conditions without retraining makes it especially attractive for sensitivity analysis, real-time prediction, design optimization, and inverse problems, where traditional numerical solvers or even physics-informed neural networks may fall short due to computational or formulation constraints.
9 Acknowledgments
This material is based on work supported by the National Science Foundation (NSF) under Grant No. CMMI 2145092. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF.
Appendix A Numerical methods to solve Equation 7
The simplest BDF method (first-order backward Euler differentiation) can be expressed in matrix form as:
$(\mathbf{I}-\Delta t\mathbf{A})\mathbf{u}^{t+1}=\mathbf{u}^{t}$ (12)
where $\mathbf{I}$ is the identity matrix, and $\mathbf{u}^{t}$ and $\mathbf{u}^{t+1}$ represent vectors of excess PWPs at the current and next time steps, respectively. The matrix $\mathbf{A}$ represents the central difference approximation for the second derivative, which arises from the spatial discretization. It has a tridiagonal matrix form:
$\mathbf{A}=\frac{C_{v}}{\Delta z^{2}}\begin{bmatrix}-2&1&0&\dots&0\\[5.0pt] 1&-2&1&\dots&0\\[5.0pt] 0&1&-2&\dots&0\\[5.0pt] \vdots&\vdots&\vdots&\ddots&1\\[5.0pt] 0&0&0&1&-2\end{bmatrix}$ (13)
At each time step, the solver implicitly solves Equation 12 for the unknown PWP distribution $\mathbf{u}^{t+1}$. In practice, the solver adaptively selects an integration order (up to fifth-order BDF) according to prescribed error tolerances. The general formulation of higher-order BDF methods is given by:
$\sum_{j=0}^{k}\alpha_{j}\mathbf{u}^{t+1-j}=\Delta t\,\mathbf{A}\mathbf{u}^{t+1}$ (14)
where $\alpha_{j}$ are method-specific coefficients determined by the selected order $k$, which denotes the number of previous time steps used in the integration scheme. SciPy uses adaptive $k$ from 1 to 5 for the optimized performance based on estimated local error and step size control.
Unlike implicit schemes, the Runge-Kutta (RK) integration method does not require solving a linear system at each step. The general form of the explicit integration step is given as:
$\mathbf{u}^{t+1}=\mathbf{u}^{t}+\Delta t\,f(t_{n},\mathbf{u}^{t}),$ (15)
where $f(\mathbf{u})=\mathbf{A}\mathbf{u}$ represents the spatially discretized PDE right-hand side.
For instance, the classical fourth-order Runge-Kutta (RK4) method computes the next step explicitly as follows:
$k_{1}$ $=f(t_{n},\mathbf{u}^{t}),$ (16)
$k_{2}$ $=f\left(t_{n}+\frac{\Delta t}{2},\mathbf{u}^{t}+\frac{\Delta t}{2}k_{1}\right),$
$k_{3}$ $=f\left(t_{n}+\frac{\Delta t}{2},\mathbf{u}^{t}+\frac{\Delta t}{2}k_{2}\right),$
$k_{4}$ $=f(t_{n}+\Delta t,\mathbf{u}^{t}+\Delta t\,k_{3}),$
$\mathbf{u}^{t+1}$ $=\mathbf{u}^{t}+\frac{\Delta t}{6}\left(k_{1}+2k_{2}+2k_{3}+k_{4}\right)$
Explicit RK methods are simpler to implement and computationally inexpensive per step, but they typically require smaller time steps compared to implicit methods. SciPy’s solve_ivp adaptively selects the time step size for improved efficiency.